Find the sum of the first 10 terms of the series: 5+15+45...
Answer by stanbon(75887)

Find the sum of the first 10 terms of the series: 5+15+45...
--------
Geometric Series with a = 5 and r = 3
-----
s(10) = a[(r^n -1)/(r-1)] = 5[(3^5-1)/(2)] = (5/2)(242) = 5*121 = 605
-------------
Cheers,
Stan H.
-------------